4. Draw the shear and moment diagrams for the compound beam. ((Lec/Tute a half) SkN/m
Answer #1

ANSWER :22 0 5 kN/m. im * free body diagram for given figure is as follows. 5 kalm Skalon tom vo. consider portion AB EMA =0 Lv8x2= 5shear calculations g sf at D = -(%) = -(-5)-5KN. State (Right of c) = -(-5)+5= 10 ka stat c. (leftofe) -(-5)+5-20 =-10KN. SESkalmi 3 B NA VB 5 kallm. him Im 10 5 + t 5 (-) с 5 SFD. 10 215 kNam 14) e B. (-) BMD 7.5kN-m M